OpenStudy (haltiamali):

Three times one number minus twice another number is -13. The first number plus three times the second number is 14. What the two numbers? Write & solve a system of equations using the method of your choice. (Substitution or elimination).

OpenStudy (anonymous):

You can call the two numbers x and y. The two equations you can create are: 3x-2y=-13 x+3y=14 Can you solve it from here or it in the solving of the system of equations that's giving you trouble?

OpenStudy (haltiamali):

I have it from here, thanks! (: Do you think you could figure this one out, too? I'm having troubles with it and it's kinda the same. /: The difference between the length and width of a rectangle is 7 centimeters. Find the dimensions of the rectangle if it's perimeter is 50 centimeters. Write & solve a system of equations by substitution.

OpenStudy (anonymous):

\(l-w=7\) \(2l+2w=50\) again...substituting \(l\) in \(2l+2w=50\) with \(l=w+7\) should make life easy :) here's a tip, whenever you have a lonely variable, solve for it and substitute

OpenStudy (anonymous):

I'm sure you can take it on from there :) @haltiamali

OpenStudy (haltiamali):

Yup! :D Thank you!

OpenStudy (anonymous):

anytime :)

OpenStudy (haltiamali):

Jamie used the elimination method to solve a system of linear equations. She discovered tha the result is 5=7. What is the solution?

OpenStudy (anonymous):

5=7 is not a solution

OpenStudy (haltiamali):

That's why I was confused..

OpenStudy (anonymous):

:)

OpenStudy (haltiamali):

So it's no solution?